OpenStudy (anonymous):

What is 5pi over 4 radians in degrees?

OpenStudy (unklerhaukus):

use pi radiens = 180°

OpenStudy (anonymous):

i still dont understand

OpenStudy (unklerhaukus):

\[\frac{5\pi}{4} [\text{rad} ]\to\frac{5}{4}\pi [\text{rad} ]\to\frac{5}{4}\times180°\]

OpenStudy (anonymous):

225 degrees right

OpenStudy (unklerhaukus):

yes,

OpenStudy (anonymous):

ok thanks man

OpenStudy (anonymous):

do you think you can help me with this What is the reference angle for 7pi over 6 ?

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):



OpenStudy (anonymous):

so is it pi over 6

OpenStudy (unklerhaukus):

yes\[\checkmark\]

OpenStudy (unklerhaukus):

what is π/6 radiens converted to degrees?

OpenStudy (anonymous):

180?

OpenStudy (unklerhaukus):

180/6 ....

OpenStudy (anonymous):

one more if you can If sin theta = 12/13 and cos theta < 0, what is tan theta ? theta is the zero with the line in it

OpenStudy (unklerhaukus):

with cos theta < 0 which quadrants must the triangle be in

OpenStudy (anonymous):

idk im not sure what youre asking